# EXAONE 3.0 7.8B Instruction Tuned Language Model

### v1

@@ 3.1.3 Coding
To evaluate coding capabilities, we measured the performance on popular benchmarks for Python code generation, focusing on relatively simple, self-contained functions. HumanEval measures functional correctness for synthesizing Python programs from docstrings, and MBPP (The Mostly Basic Programming Problems) measures models’ ability to synthesize short Python programs. As shown in Table  9, EXAONE 3.0 7.8B instruction-tuned model’s performance in HumanEval stands out compared to other models, and it also shows competitive performance in MBPP. Consequently, the average scores indicate that our model demonstrated superior coding capability compared to other models.
@@
Table 9: Evaluation results of EXAONE 3.0 7.8B instruction-tuned model on two Python code generation benchmarks. Our model excels on the HumanEval benchmark with the highest score, while demonstrating competitive results on the MBPP benchmark compared to baseline models. As a result, our model achieved top performance on average across these two code generation benchmarks. The performance was measured using the pass@1 score, with a zero-shot prompt for HumanEval and a 3-shot prompt for MBPP.
@@ 8.2 Benchmarks
MBPP pass@1 (3-shot)
@@ 8.3 Model License
EXAONE AI Model License Agreement 1.0 - NC
@@ 2. License Grant
a. Access, download, install, and use the Model solely for internal research purposes. This includes evaluation, testing, academic research, experimentation, and participation in competitions, provided that such participation is in a non-commercial context. Notwithstanding Section 3.1, the Licensee may only provide the Model or Derivatives for a competition if no license is granted to the competition organizer or any third party.
b. Publicly disclose research results and findings derived from the use of the Model or Derivatives, including publishing papers or presentations, but may not distribute, share, or otherwise make the Model or any Derivatives available to third parties. When publishing or presenting, only a minimal amount of Output necessary for disclosing research results should be made public.
c. Modify the Model and create Derivatives based on the Model, provided that such modifications and Derivatives are used exclusively for internal research purposes. The Licensee may conduct experiments, perform analyses, and apply custom modifications to the Model to explore its capabilities and performance under various scenarios.
@@ 3. Restrictions
3.1 Sublicensing and Distribution: The Licensee shall not sublicense, sell, rent, lease, distribute, disclose, or otherwise transfer the Model, any Derivatives, or any Output to any third party, except as permitted in Section 2.1. This prohibition extends to making the Model available as part of a service, product, or solution offered to others, whether for free or for compensation.
@@ 4. Ownership
4.2 Output: All rights, title, and interest in and to the Output generated by the Model, whether in its original form or modified, are and shall remain the exclusive property of the Licensor. The Licensee shall not claim ownership of the Output except as expressly provided in this Agreement. The Licensee may use the Output solely for the purposes permitted under this Agreement and shall not exploit the Output for unauthorized or commercial purposes.

### v2

@@ 3.1.3 Coding
To evaluate coding capabilities, we measured the performance on popular benchmarks for Python code generation, focusing on relatively simple, self-contained functions. HumanEval measures functional correctness for synthesizing Python programs from docstrings, and MBPP (The Mostly Basic Programming Problems) measures models’ ability to synthesize short Python programs. As shown in Table  3, EXAONE 3.0 7.8B instruction-tuned model’s performance in HumanEval stands out compared to other models, and it also shows competitive performance in MBPP. Consequently, the average scores indicate that our model demonstrated superior coding capability compared to other models.
@@
Table 9: Evaluation results of EXAONE 3.0 7.8B instruction-tuned model on two Python code generation benchmarks. Our model excels on the HumanEval benchmark with the highest score, while demonstrating competitive results on the MBPP benchmark compared to baseline models. As a result, our model achieved top performance on average across these two code generation benchmarks. The performance was measured using the pass@1 score, with a zero-shot prompt for both HumanEval and MBPP  [^3] We assessed the code generation performance of language models using the default settings of the BigCode evaluation harness environment [8]. All models were evaluated with zero-shot prompts, without their own chat templates..
@@ 8.2 Benchmarks
MBPP pass@1 (0-shot)
@@ 8.3 Model License
EXAONE AI Model License Agreement 1.1 - NC
@@ 2. License Grant
a. Access, download, install, and use the Model solely for research purposes. This includes evaluation, testing, academic research, experimentation, and participation in competitions, provided that such participation is in a non-commercial context. Notwithstanding Section 3.1, the Licensee may only provide the Model or Derivatives for a competition if no commercial license is granted to the competition organizer or any third party.
b. Publicly disclose research results and findings derived from the use of the Model or Derivatives, including publishing papers or presentations.
c. Modify the Model and create Derivatives based on the Model, provided that such modifications and Derivatives are used exclusively for research purposes. The Licensee may conduct experiments, perform analyses, and apply custom modifications to the Model to explore its capabilities and performance under various scenarios. If the Model is modified, the modified Model must include “EXAONE” at the beginning of its name.
d. Distribute the Model and Derivatives in each case with a copy of this Agreement.
@@ 4. Ownership
4.2 Output: All rights, title, and interest in and to the Output generated by the Model and Derivatives whether in its original form or modified, are and shall remain the exclusive property of the Licensor. Licensee may use, modify, and distribute the Output and its derivatives for research purpose. The Licensee shall not claim ownership of the Output except as expressly provided in this Agreement. The Licensee may use the Output solely for the purposes permitted under this Agreement and shall not exploit the Output for unauthorized or commercial purposes.